OpenStudy (anonymous):

One person can do a certain job in ten minutes, and another person can do the same job in fifteen minutes. How many minutes will they take to do the job together?

hero (hero):

Hint: Pretend that the first person is a man and the second person is a woman. Then: \(\large\frac{M \times W}{M + W}\)

OpenStudy (anonymous):

Okay...

hero (hero):

Thus: \(\large\frac{10 \times 15}{10 + 15}\)

OpenStudy (anonymous):

150/25=6

hero (hero):

So they can both do the job in 6 minutes.

OpenStudy (anonymous):

Yes!!! Thank you!